# BiT-MamSleep: Bidirectional Temporal Mamba for EEG Sleep Staging

### v1

@@ Abstract
In this paper, we address the challenges in automatic sleep stage classification, particularly the high computational cost, inadequate modeling of bidirectional temporal dependencies, and class imbalance issues faced by Transformer-based models. To address these limitations, we propose BiT-MamSleep, a novel architecture that integrates the Triple-Resolution CNN (TRCNN) for efficient multi-scale feature extraction with the Bidirectional Mamba (BiMamba) mechanism, which models both short- and long-term temporal dependencies through bidirectional processing of EEG data. Additionally, BiT-MamSleep incorporates an Adaptive Feature Recalibration (AFR) module and a temporal enhancement block to dynamically refine feature importance, optimizing classification accuracy without increasing computational complexity. To further improve robustness, we apply optimization techniques such as Focal Loss and SMOTE to mitigate class imbalance. Extensive experiments on four public datasets demonstrate that BiT-MamSleep significantly outperforms state-of-the-art methods, particularly in handling long EEG sequences and addressing class imbalance, leading to more accurate and scalable sleep stage classification.
@@ I Introduction
Sleep is a vital physiological process that plays a crucial role in maintaining health and well-being. Poor sleep is linked to various serious health conditions, including cardiovascular diseases, cognitive impairments, and mental health disorders [1, 2]. Understanding and classifying sleep stages is essential for diagnosing sleep disorders, which affect a significant portion of the population. Electroencephalogram (EEG) signals, which capture brain activity during sleep, serve as a critical tool for analyzing these stages, as they provide insight into the physiological and neurological patterns associated with different sleep phases. EEG-based sleep stage classification is thus crucial for enabling accurate diagnoses and guiding effective treatments.
@@
The high computational cost of Transformers poses a significant obstacle when processing long sleep EEG sequences. Sleep EEG data, typically recorded at high sampling rates over extended durations, consist of a large number of time steps, resulting in highly complex datasets. Transformers, due to their quadratic increase in computational complexity as sequence length grows, struggle to efficiently process such data. This computational inefficiency poses a significant limitation, particularly in real-world sleep stage classification tasks where handling long EEG sequences with scalability is crucial. Therefore, it is crucial to develop models with reduced computational overhead and enhanced scalability for processing complex EEG data.
@@
Class imbalance further complicates sleep stage classification, highlighting another limitation of Transformer-based models. Sleep stages like REM occur far less frequently than others, such as NREM, leading to skewed predictions toward the majority classes. Transformers, lacking mechanisms to address this imbalance, tend to underperform when classifying minority stages. This imbalance negatively impacts overall classification accuracy, particularly in detecting underrepresented sleep stages. Effective handling of class imbalance is essential to improve the performance of models in sleep stage classification tasks, making it a key area for optimization in future approaches.
@@
To address these challenges, we integrate the feature extraction capabilities of the Triple-Resolution CNN (TRCNN) with the Bidirectional State Space Model (BSSM). The TRCNN efficiently captures multi-scale frequency features essential for accurate sleep stage identification, while the BiMamba compensates for Transformer limitations by modeling both short- asnd long-term temporal dependencies through bidirectional EEG data processing. We further enhance model performance by introducing an Adaptive Feature Recalibration (AFR) module, which dynamically refine feature importance, optimizing classification accuracy without increasing computational complexity. Additionally, we propose a series of data optimization techniques to mitigate class imbalance and improve model stability during training, ensuring robust generalization across diverse sleep datasets.
@@
    We propose BiT-MamSleep, a novel architecture that integrates the Triple-Resolution CNN (TRCNN) with the BiMamba mechanism. This design efficiently captures multi-scale spatial and temporal features from EEG data while maintaining linear computational complexity, allowing for scalable and efficient processing of long EEG sequences.
@@
    Bidirectional Temporal Modeling: The BiMamba mechanism utilizes Bidirectional State Space Models (BSSM) which captures both past and future temporal dependencies, improving the accuracy of detecting transitions
@@ II-A Early approaches of automatic sleep stage classification
Traditional sleep stage classification methods relied on expert manual annotation of EEG signals, which was time-consuming and labor-intensive. As sleep studies expanded, manual annotation became impractical, prompting the development of automated systems using machine learning techniques, such as support vector machines (SVMs) and other models [39]. While these methods reduced human involvement by automating feature extraction, they depended on hand-crafted features, which limited their generalizability across different datasets, subjects, and recording conditions. Variations in EEG recordings between individuals posed challenges to model robustness, and these models often struggled to maintain accuracy when confronted with new, unseen data. As more diverse EEG data became available, the limitations of feature engineering became evident, leading to the adoption of more flexible and powerful approaches such as deep learning.
@@ II-B Deep Learning Models: CNNs and RNNs
The introduction of deep learning marked a significant shift in sleep stage classification, particularly through the use of convolutional neural networks (CNNs) [11, 12, 13]. CNNs demonstrated effectiveness in capturing local spatial features from raw EEG signals, eliminating the need for manual feature engineering. However, CNNs’ fixed receptive fields limit their ability to model long-term dependencies, which are crucial for detecting transitions between sleep stages like REM and NREM. Recurrent neural networks (RNNs), designed to handle sequential data, offered an alternative for modeling temporal dependencies [14, 15, 16]. Despite their advantages, RNNs faced the well-documented vanishing and exploding gradient problems when dealing with long EEG sequences. These issues hindered RNNs’ effectiveness in accurately capturing long-term dependencies, revealing the need for a model that could efficiently handle both short- and long-term dependencies in sleep data.
@@ II-D State Space Models (SSMs) and the Mamba Mechanism
State Space Models (SSMs) [25, 26, 27] provide an efficient alternative to Transformers by capturing complex temporal dependencies with linear computational complexity, making them suitable for processing long-duration EEG sequences. Mamba further optimizes SSMs through a hardware-aware design and dynamic time-step selection strategy, improving computational efficiency and ensuring the model focuses on the most relevant EEG segments [24]. However, both SSMs and Mamba still struggle with capturing bidirectional dependencies, which are critical for understanding sleep stage transitions
@@ III-A Feature Extraction
The Feature Extraction Module is designed to effectively extract multi-scale and multi-frequency features from raw EEG signals, thereby enhancing the accuracy of sleep stage classification. Sleep stage classification relies on extracting frequency-specific features from EEG signals, as different sleep stages are characterized by distinct frequency bands [16]. To achieve this goal, the Feature Extraction Module consists of two main components. First, the Multi-Resolution Convolutional Neural Network (MRCNN) is employed to capture rich temporal and frequency-domain features at various scales. Then, the Adaptive Feature Recalibration (AFR) module dynamically adjusts and optimizes these features by learning the interdependencies between them. Below, we provide a detailed explanation of each of these components, as illustrated in Fig. 2.
@@ III-A1 Triple-Resolution Convolutional Neural Network
To effectively extract features from raw EEG signals, we developed a Triple-Resolution Convolutional Neural Network (TRCNN) with three distinct branches of convolutional layers. Each branch employs a unique kernel size to capture specific temporal and frequency characteristics of the EEG signals. The use of three branches comprehensively addresses the non-stationary nature of EEG signals, where different sleep stages exhibit distinct frequency bands, such as delta waves in deep sleep (N3) and alpha waves in light sleep (N1). Compared to dual-branch designs, this multi-resolution approach enables our model to capture a broader spectrum of frequency-specific information, thereby enhancing the feature representation for sleep stage classification.
@@ III-A2 Adaptive Feature Modules (AFMs)
To enhance feature extraction, we designed Adaptive Feature Modules, decomposing the original Adaptive Feature Recalibration (AFR) into three submodules: Feature Squeeze (FSq), Feature Excitation (FE), and Feature Scale (FSs). This modular approach incrementally improves feature representation
@@
  • Feature Excitation (FE): The Feature Excitation module utilizes a two-layer fully connected network, similar to the SE layer [33], to learn dependencies between different feature channels and adaptively adjust their importance to enhance the feature representation. The first layer reduces dimensionality using ReLU, and the second restores it using Sigmoid, allowing the model to focus on discriminative features, thereby improving classification accuracy.
  • Feature Scale (FSs): The Feature Scale module combines the excited features from Feature Excitation to better capture inter-scale feature relationships. Through a channel-wise scaling operation, each channel of the input features is reweighted, emphasizing important features and suppressing irrelevant ones to enhance the final feature representation. A shortcut connection is added to retain original information and ensure stability [34].
@@ IV-A1 Sleep-EDF-20 and Sleep-EDF-78
The Sleep-EDF-20 and Sleep-EDF-78 datasets were obtained from PhysioBank [37]. Sleep-EDF-20 contains data from 20 healthy subjects aged 25 to 34, recorded during two consecutive nights as part of the Sleep Cassette (SC) study, which focused on the effects of aging on sleep. Sleep-EDF-78, an expanded version, includes 78 recordings from both the SC and Sleep Telemetry (ST) studies. While SC investigates aging effects, the ST study examines the impact of temazepam on sleep in 22 Caucasian participants. Each PSG file contains two EEG channels (Fpz-Cz, Pz-Oz), one EOG channel, and one chin EMG channel, all sampled at 100 Hz. For consistency with previous studies [38], [39], [40], we used only the Fpz-Cz channel as input in our experiments.
@@ IV-A2 Sleep Heart Health Study (SHHS)
The Sleep Heart Health Study (SHHS) dataset, a multi-center cohort study, focuses on the cardiovascular and other health consequences of sleep-disordered breathing [41]. From the original 6,441 participants, we selected 329 subjects with regular sleep patterns (Apnea-Hypopnea Index, AHI ¡ 5) to minimize the influence of comorbidities such as lung and cardiovascular diseases [42]. For our experiments, we used the C4-A1 EEG channel, originally sampled at 125 Hz and downsampled to 100 Hz for consistency with other datasets. The recordings were segmented into 30-second epochs in accordance with AASM guidelines.
@@ IV-A3 ISRUC-S3 Dataset
The ISRUC-S3 dataset represents a crucial advancement in the field of sleep stage classification, offering an extensive collection of polysomnographic (PSG) recordings from individuals with potential sleep disorders. Unlike traditional datasets, ISRUC-S3 provides high-resolution, clinically annotated data across multiple channels, including EEG, EOG, and EMG, recorded overnight and segmented into the standard 30-second epochs. The annotations, meticulously labeled by certified sleep technicians following the American Academy of Sleep Medicine (AASM) guidelines, encompass the five primary sleep stages (W, N1, N2, N3, and REM), ensuring robust support for multi-stage classification tasks [43].This dataset was sampled at 200 Hz, and we selected the C3-A2 channel as input for the model. Although this differs from the channels used in the other datasets, the choice of C3-A2 for ISRUC-S3 aligns with previous studies and maximizes the data quality available in this dataset.
@@ IV-B Baselines and Experimental Setup
To address class imbalance, we employed Focal Loss, particularly focusing on the underrepresented sleep stages (e.g., N1 and REM). We incorporated an early stopping strategy that halted training if the validation accuracy did not improve after 30 epochs. If there was no improvement for 15 consecutive epochs, we switched from the ReduceLROnPlateau scheduler to a CyclicLR scheduler. This induced periodic learning rate oscillations to help the model escape local minima.
Further hyperparameter details, including the TCE module configuration and dropout rates, can be found in our publicly available code at GitHub repository link.

### v2

@@ Abstract
In this paper, we address the challenges in automatic sleep stage classification, particularly the high computational cost, inadequate modeling of bidirectional temporal dependencies, and class imbalance issues faced by Transformer-based models. To address these limitations, we propose BiT-MamSleep, a novel architecture that integrates the Triple-Resolution Convolutional Neural Network (TRCNN) for efficient multi-scale feature extraction with the Bidirectional Mamba (BiMamba) mechanism, which models both short- and long-term temporal dependencies through bidirectional processing of electroencephalogram (EEG) data. Additionally, BiT-MamSleep incorporates an Adaptive Feature Recalibration (AFR) module and a temporal enhancement block to dynamically refine feature importance, optimizing classification accuracy without increasing computational complexity. To further improve robustness, we apply optimization techniques such as Focal Loss and SMOTE to mitigate class imbalance. Extensive experiments on four public datasets demonstrate that BiT-MamSleep significantly outperforms state-of-the-art methods, particularly in handling long EEG sequences and addressing class imbalance, leading to more accurate and scalable sleep stage classification.
@@ I Introduction
Sleep is a vital physiological process that plays a crucial role in maintaining health and well-being. Poor sleep is linked to various serious health conditions, including cardiovascular diseases, cognitive impairments, and mental health disorders [1, 2]. Understanding and classifying sleep stages is essential for diagnosing sleep disorders, which affect a significant portion of the population [3, 4]. Electroencephalogram (EEG) signals, which capture brain activity during sleep, serve as a critical tool for analyzing these stages, as they provide insight into the physiological and neurological patterns associated with different sleep phases. EEG-based sleep stage classification is thus crucial for enabling accurate diagnoses and guiding effective treatments.
@@
The high computational cost of Transformers poses a significant obstacle when processing long sleep EEG sequences [9]. Sleep EEG data, typically recorded at high sampling rates over extended durations, consist of a large number of time steps, resulting in highly complex datasets. Transformers, due to their quadratic increase in computational complexity as sequence length grows, struggle to efficiently process such data. This computational inefficiency poses a significant limitation, particularly in real-world sleep stage classification tasks where handling long EEG sequences with scalability is crucial. Therefore, it is crucial to develop models with reduced computational overhead and enhanced scalability for processing complex EEG data [10, 11, 12, 13].
@@
Class imbalance further complicates sleep stage classification, highlighting another limitation of Transformer-based models [14, 15, 16]. Sleep stages like REM occur far less frequently than others, such as NREM, leading to skewed predictions toward the majority classes. Transformers, lacking mechanisms to address this imbalance, tend to underperform when classifying minority stages. This imbalance negatively impacts overall classification accuracy, particularly in detecting underrepresented sleep stages. Effective handling of class imbalance is essential to improve the performance of models in sleep stage classification tasks, making it a key area for optimization in future approaches.
@@
To address these challenges, we integrate the feature extraction capabilities of the Triple-Resolution Convolutional Neural Network (TRCNN) with the Bidirectional State Space Model (BSSM). The TRCNN efficiently captures multi-scale frequency features essential for accurate sleep stage identification, while the Bidirectional Mamba (BiMamba) compensates for Transformer limitations by modeling both short- asnd long-term temporal dependencies through bidirectional EEG data processing. We further enhance model performance by introducing an Adaptive Feature Recalibration (AFR) module, which dynamically refine feature importance, optimizing classification accuracy without increasing computational complexity. Additionally, we propose a series of data optimization techniques to mitigate class imbalance and improve model stability during training, ensuring robust generalization across diverse sleep datasets.
@@
    We propose BiT-MamSleep, a novel architecture that integrates the TRCNN with the BiMamba mechanism. This design efficiently captures multi-scale spatial and temporal features from EEG data while maintaining linear computational complexity, allowing for scalable and efficient processing of long EEG sequences.
@@
    Bidirectional Temporal Modeling: The BiMamba mechanism utilizes BSSM which captures both past and future temporal dependencies, improving the accuracy of detecting transitions
@@ II-A Early approaches of automatic sleep stage classification
Traditional sleep stage classification methods relied on expert manual annotation of EEG signals, which was time-consuming and labor-intensive. As sleep studies expanded, manual annotation became impractical, prompting the development of automated systems using machine learning techniques, such as SVMs and other models [44]. While these methods reduced human involvement by automating feature extraction, they depended on hand-crafted features, which limited their generalizability across different datasets, subjects, and recording conditions. Variations in EEG recordings between individuals posed challenges to model robustness, and these models often struggled to maintain accuracy when confronted with new, unseen data. As more diverse EEG data became available, the limitations of feature engineering became evident, leading to the adoption of more flexible and powerful approaches such as deep learning.
@@ II-B Deep Learning Models: CNNs and RNNs
The introduction of deep learning marked a significant shift in sleep stage classification, particularly through the use of CNNs [17, 18]. CNNs demonstrated effectiveness in capturing local spatial features from raw EEG signals, eliminating the need for manual feature engineering. However, CNNs’ fixed receptive fields limit their ability to model long-term dependencies, which are crucial for detecting transitions between sleep stages like REM and NREM. RNNs, designed to handle sequential data, offered an alternative for modeling temporal dependencies [19, 20, 21]. Despite their advantages, RNNs faced the well-documented vanishing and exploding gradient problems when dealing with long EEG sequences. These issues hindered RNNs’ effectiveness in accurately capturing long-term dependencies, revealing the need for a model that could efficiently handle both short- and long-term dependencies in sleep data.
@@ II-D State Space Models (SSMs) and the Mamba Mechanism
State Space Models (SSMs) [30, 31, 32] provide an efficient alternative to Transformers by capturing complex temporal dependencies with linear computational complexity, making them suitable for processing long-duration EEG sequences. Mamba further optimizes SSMs through a hardware-aware design and dynamic time-step selection strategy, improving computational efficiency and ensuring the model focuses on the most relevant EEG segments [29]. However, both SSMs and Mamba still struggle with capturing bidirectional dependencies, which are critical for understanding sleep stage transitions.
@@ III-A Feature Extraction
The Feature Extraction Module is designed to effectively extract multi-scale and multi-frequency features from raw EEG signals, thereby enhancing the accuracy of sleep stage classification. Sleep stage classification relies on extracting frequency-specific features from EEG signals, as different sleep stages are characterized by distinct frequency bands [21]. To achieve this goal, the Feature Extraction Module consists of two main components. First, the Multi-Resolution Convolutional Neural Network (MRCNN) is employed to capture rich temporal and frequency-domain features at various scales. Then, the AFR module dynamically adjusts and optimizes these features by learning the interdependencies between them. Below, we provide a detailed explanation of each of these components, as illustrated in Fig. 2.
@@ III-A1 Triple-Resolution Convolutional Neural Network
To effectively extract features from raw EEG signals, we developed a TRCNN with three distinct branches of convolutional layers. Each branch employs a unique kernel size to capture specific temporal and frequency characteristics of the EEG signals. The use of three branches comprehensively addresses the non-stationary nature of EEG signals, where different sleep stages exhibit distinct frequency bands, such as delta waves in deep sleep (N3) and alpha waves in light sleep (N1). Compared to dual-branch designs, this multi-resolution approach enables our model to capture a broader spectrum of frequency-specific information, thereby enhancing the feature representation for sleep stage classification.
@@ III-A2 Adaptive Feature Modules (AFMs)
To enhance feature extraction, we designed Adaptive Feature Modules, decomposing the original AFR into three submodules: Feature Squeeze (FSq), Feature Excitation (FE), and Feature Scale (FSs). This modular approach incrementally improves feature representation
@@
  • FE: The Feature Excitation module utilizes a two-layer fully connected network, similar to the SE layer [38], to learn dependencies between different feature channels and adaptively adjust their importance to enhance the feature representation. The first layer reduces dimensionality using ReLU, and the second restores it using Sigmoid, allowing the model to focus on discriminative features, thereby improving classification accuracy.
  • FSs: The Feature Scale module combines the excited features from Feature Excitation to better capture inter-scale feature relationships. Through a channel-wise scaling operation, each channel of the input features is reweighted, emphasizing important features and suppressing irrelevant ones to enhance the final feature representation. A shortcut connection is added to retain original information and ensure stability [39].
@@ IV-A1 Sleep-EDF-20 and Sleep-EDF-78
The Sleep-EDF-20 and Sleep-EDF-78 datasets were obtained from PhysioBank [42]. Sleep-EDF-20 contains data from 20 healthy subjects aged 25 to 34, recorded during two consecutive nights as part of the Sleep Cassette (SC) study, which focused on the effects of aging on sleep. Sleep-EDF-78, an expanded version, includes 78 recordings from both the SC and Sleep Telemetry (ST) studies. While SC investigates aging effects, the ST study examines the impact of temazepam on sleep in 22 Caucasian participants. Each polysomnographic (PSG) file contains two EEG channels (Fpz-Cz, Pz-Oz), one EOG channel, and one chin EMG channel, all sampled at 100 Hz. For consistency with previous studies [43], [44], [45], we used only the Fpz-Cz channel as input in our experiments.
@@ IV-A2 Sleep Heart Health Study (SHHS)
The SHHS dataset, a multi-center cohort study, focuses on the cardiovascular and other health consequences of sleep-disordered breathing [46]. From the original 6,441 participants, we selected 329 subjects with regular sleep patterns to minimize the influence of comorbidities such as lung and cardiovascular diseases [47]. For our experiments, we used the C4-A1 EEG channel, originally sampled at 125 Hz and downsampled to 100 Hz for consistency with other datasets. The recordings were segmented into 30-second epochs in accordance with American Academy of Sleep Medicine (AASM) guidelines.
@@ IV-A3 ISRUC-S3 Dataset
The ISRUC-S3 dataset represents a crucial advancement in the field of sleep stage classification, offering an extensive collection of PSG recordings from individuals with potential sleep disorders. Unlike traditional datasets, ISRUC-S3 provides high-resolution, clinically annotated data across multiple channels, including EEG, EOG, and EMG, recorded overnight and segmented into the standard 30-second epochs. The annotations, meticulously labeled by certified sleep technicians following the AASM guidelines, encompass the five primary sleep stages (W, N1, N2, N3, and REM), ensuring robust support for multi-stage classification tasks [48].This dataset was sampled at 200 Hz, and we selected the C3-A2 channel as input for the model. Although this differs from the channels used in the other datasets, the choice of C3-A2 for ISRUC-S3 aligns with previous studies and maximizes the data quality available in this dataset.
@@ IV-B Baselines and Experimental Setup
To address class imbalance, we employed Focal Loss, particularly focusing on the underrepresented sleep stages (e.g., N1 and REM). We incorporated an early stopping strategy that halted training if the validation accuracy did not improve after 30 epochs. If there was no improvement for 15 consecutive epochs, we switched from the RRP scheduler to a CyclicLR scheduler. This induced periodic learning rate oscillations to help the model escape local minima.
@@ IV-C Experiment Results
The BiT-MamSleep was extensively evaluated across four well-established sleep datasets, Sleep-EDF-20, Sleep-EDF-78, SHHS, and ISRUC-S3, to rigorously assess its capacity for accurate, robust, and balanced sleep stage classification. Each dataset presents unique challenges, from class imbalance to diverse EEG patterns, enabling a comprehensive comparison with aforementioned baseline models: Random Forest, DeepSleepNet, ResNetLSTM, and AttnSleep. The results reveal that BiT-MamSleep significantly outperforms these models in both overall accuracy and in handling critical classification challenges, including long-range temporal dependencies, multi-scale feature extraction, and imbalanced sleep stage data.
On the Sleep-EDF-20 dataset, BiT-MamSleep achieved an accuracy of 85.18%, outperforming established models by at least 2% and showcasing superior robustness with a macro-averaged F1-score of 78.51% and a Cohen’s Kappa of 79.59%. This dataset’s imbalanced distribution, where common stages like N2 dominate over less frequent ones like N1 and REM, which pose substantial challenges to baseline models. For instance, DeepSleepNet, which combines CNN and LSTM layers, exhibited limitations in capturing the temporal transitions necessary for accurately detecting minority stages. Similarly, ResNetLSTM, though designed to capture temporal patterns, struggled with the dataset’s long EEG sequences due to gradient-related issues inherent in recurrent architectures, often resulting in reduced accuracy on rare stages. AttnSleep, which employs attention to capture temporal dependencies, performed better overall but is constrained by its unidirectional processing, limiting its ability to fully capture bidirectional context in stage transitions.
In contrast, BiT-MamSleep addresses these limitations with its BiMamba mechanism, which enables the model to learn both forward and backward dependencies, crucial for accurately modeling sleep stage transitions over extended EEG sequences. This dual-directional processing proved particularly advantageous in identifying stages like N1 and REM, where subtle temporal dependencies are essential. Furthermore, BiT-MamSleep’s adaptive feature recalibration dynamically adjusts the relevance of specific features across stages, allowing it to effectively classify even the minority stages, thereby enhancing the model’s balanced performance across all classes.
On the Sleep-EDF-78 dataset, BiT-MamSleep maintained its edge, achieving an accuracy of 80.21%. This dataset, characterized by a broader subject base and greater EEG variability, challenged baseline models to generalize effectively. While AttnSleep’s self-attention mechanism improved its performance over recurrent models by capturing longer dependencies, its single-pass processing lacked the dual temporal context provided by BiT-MamSleep’s bidirectional mechanism. BiT-MamSleep’s TRCNN played a crucial role here, enabling multi-resolution feature extraction that effectively captures both high-frequency details and long-range dependencies in EEG signals. This multi-resolution approach allowed BiT-MamSleep to have good performance in the classification of minority stages, such as N1, where previous models showed weaker performance. The adaptability of BiT-MamSleep’s architecture to capture both spatial and temporal characteristics ensured consistent gains over prior models.
BiT-MamSleep further demonstrated its adaptability on the SHHS dataset, achieving an accuracy of 84.38%. This dataset, which includes data from subjects with sleep-disordered breathing, introduces unique EEG patterns that increase classification complexity. Random Forest, as a non-deep-learning baseline, struggled with these challenges due to its reliance on hand-crafted features, which do not capture the intricate temporal patterns inherent in EEG. Both DeepSleepNet and ResNetLSTM encountered limitations in handling the distinct temporal dependencies introduced by disordered breathing events, as their sequential modeling approaches lack the bidirectional perspective essential for comprehensive temporal coverage. BiT-MamSleep’s bidirectional mechanism again proved instrumental, allowing the model to process both historical and future context in parallel, crucial for capturing the physiological nuances that distinguish each sleep stage. Additionally, the model’s balanced performance, reflected in a high macro-averaged G-mean, confirms its ability to classify each stage accurately, avoiding the skew often seen in models not equipped to handle imbalanced classes.
On the ISRUC-S3 dataset, which provides high-resolution polysomnographic data from patients with potential sleep disorders, BiT-MamSleep achieved 75.22% accuracy, outperforming all baselines. This dataset’s complex nature, with fine-grained annotations across EEG, EOG, and EMG channels, necessitates multi-resolution processing to identify sleep stages accurately. AttnSleep’s single-resolution attention model lacked the capacity to capture the diverse spectral and temporal features, particularly in stages like REM that require both long-range dependencies and high-frequency detail. BiT-MamSleep’s TRCNN architecture, by employing small, medium, and large convolutional kernels, effectively captured a wide spectrum of EEG frequencies, ensuring reliable classification even in high-resolution and diverse data. The adaptive feature recalibration module further strengthened BiT-MamSleep’s capacity to adapt to the variability of disordered sleep data, dynamically prioritizing stage-specific features, thus ensuring a balanced performance across the five main stages.
Across all datasets, BiT-MamSleep consistently demonstrated its superiority in addressing the limitations of traditional and recent deep learning models through the following advancements:
  • Bidirectional Temporal Modeling: The BiMamba mechanism’s capacity to capture both forward and backward dependencies effectively overcomes the unidirectional limitations of models like AttnSleep and DeepSleepNet, which struggle with stage transitions and lack temporal robustness.
  • Multi-Resolution Feature Extraction: The TRCNN architecture’s ability to capture frequency-specific features at various temporal scales enables BiT-MamSleep to handle the non-stationary nature of EEG data, providing it an advantage over both Random Forest and single-resolution deep learning models.
  • Dynamic Feature Recalibration: By adapting feature importance dynamically, BiT-MamSleep enhances its performance on minority classes, addressing class imbalance challenges more effectively than models like ResNetLSTM, which lack this flexibility.
  • Effective Class Imbalance Handling: With the use of focal loss and SMOTE-based resampling, BiT-MamSleep achieves balanced sensitivity across all stages, significantly improving accuracy for minority classes like N1 and REM, where traditional models often fall short.
IV-D Ablation Study